# AnyTop: Character Animation Diffusion with Any Topology

### v1

@@ Abstract.
Generating motion for arbitrary skeletons is a longstanding challenge in computer graphics, remaining largely unexplored due to the scarcity of diverse datasets and the irregular nature of the data. In this work, we introduce AnyTop, a diffusion model that generates motions for diverse characters with distinct motion dynamics, using only their skeletal structure as input. Our work features a transformer-based denoising network, tailored for arbitrary skeleton learning, integrating topology information into the traditional attention mechanism. Additionally, by incorporating textual joint descriptions into the latent feature representation, AnyTop learns semantic correspondences between joints across diverse skeletons. Our evaluation demonstrates that AnyTop generalizes well, even with as few as three training examples per topology, and can produce motions for unseen skeletons as well. Furthermore, our model’s latent space is highly informative, enabling downstream tasks such as joint correspondence, temporal segmentation and motion editing. Our webpage, https://anytop2025.github.io/Anytop-page, includes links to videos and code.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
Figure 1. AnyTop generates motions for diverse characters with distinct motion dynamics, using only their skeletal structure as input.
@@ 1. Introduction
2. Related Work
[…]
Skeletal variability in generative motion models
@@ 2. Related Work
We refer to four types of skeletal variability (table 1). The naming draws from terminology in the graph domain, hence we interchangeably use the terms joint and vertex, as well as edge and bone. A single skeleton type refers to identical skeletons — that is, skeletons with the same vertices, connectivity, and edge lengths. Isomorphic skeletons correspond to isomorphic graphs, sharing vertices and edges but potentially differing in edge proportions. Homeomorphic skeletons may vary in structure, yet correspond to homeomorphic graphs, i.e., use topologies obtained from the same primal graph by subdivision of edges. Specifically, homeomorphic skeletons share the same number of kinematic chains and end-effectors. Finally, non-homeomorphic skeletons vary in their structure and have no common primal graph.
@@
Most motion generative methods focus on a single skeletal structure (Raab et al., 2023; Petrovich et al., 2021; Karunratanakul et al., 2023). Others train on isomorphic skeletons (Zhang et al., 2023b; Villegas et al., 2021), including works that use the SMPL (Loper et al., 2015) body model (Tripathi et al., 2025; Jang et al., 2024; Petrovich et al., 2022a) and SMAL (Zuffi et al., 2017) body model or its derivatives (Yang et al., 2023; Rueegg et al., 2023). A smaller portion of generative works support homeomorphic skeletons (Lee et al., 2023; Zhang et al., 2024d; Cao and Yang, 2024; Studer et al., 2024; Ponton et al., 2024). Among these works, some (Aberman et al., 2020) require a designated encoder and decoder per skeleton, and some (Zhang et al., 2024b) offer a unified framework for all skeletons.
@@
Only a handful of works can handle non-homeomorphic skeletons. Martinelli et al. (2024) performs motion retargeting by learning a shared manifold for all skeletons, and decoding it to motions using learned skeleton-specific tokens. The learned tokens capture the skeletal information of characters in the dataset, limiting the model’s ability to generalize to skeletons unseen during training. Its results are shown exclusively on bipeds, leaving the applicability to other character families (e.g., quadrupeds, millipedes) unexplored. WalkTheDog (Li et al., 2024) uses a latent space that encodes motion phases and accommodates non-generative motion matching.
@@
A different class of generative models bypasses the handling of the skeletal structure by generating motion directly from point clouds (Mo et al., 2025), shape-handles (Zhang et al., 2023a) or meshes (Song et al., 2023; Ye et al., 2024; Muralikrishnan et al., 2024; Zhang et al., 2024c). These works demonstrate great flexibility in target character structure, but overlook the advantage of skeletons, which are more compact and semantically meaningful, easier to manipulate via rig-based animation, and compatible with physics engines (Tevet et al., 2024) and inverse kinematics systems. Some works (Wang et al., 2024) perform automatic rigging after the generation, but automatic rigging often necessitates manual adjustments.
@@ 3.1. Preliminaries
We represent motion as a 3D tensor $X\in\mathbb{R}^{N\times J\times D}$, where $N$ and $J$ are the maximum number of frames and joints across all motions in the dataset, and D is the number of motion features per joint. As motions vary in duration and skeletal structure, we pad the original number of frames and joints of each motion to match the maximum values $N$ and $J$, respectively. We adopt a redundant representation, where each joint $j$ (except the root) consists of its root-relative position $p_{j}\in\mathbb{R}^{3}$, 6D joint rotation $r_{j}\in\mathbb{R}^{6}$ (Zhou et al., 2018), linear velocity $v_{j}\in\mathbb{R}^{3}$, and foot contact label $fc_{j}\in\{0,1\}$. Altogether a joint is represented by $\{p_{j},r_{j},v_{j},fc_{j}\}\in\mathbb{R}^{13}$, hence $D=13$. For the root joint, features include its rotational velocity, linear velocity and height, which are concatenated and zero-padded to match the size $D$. Our representation is inspired by Guo et al. (2022); however, our approach maintains features at the joint level by representing each joint as a separate tensor, resulting in $J$ tokens per frame. In contrast, Guo et al. concatenate features from all joints into one tensor, resulting in a single token per frame.
@@ 3.2. Architecture
This block incorporates semantic information from the rest-pose $\mathcal{P}_{\mathcal{S}}$ and the joint descriptions $\mathcal{N}_{\mathcal{S}}$, into the noised sample $X_{t}$. It projects $\mathcal{P}_{\mathcal{S}}$ to feature lentgh $F$ and concatenates it with $X_{t}$ along the temporal axis, effectively making it frame 0. The joint descriptions $\mathcal{N}_{\mathcal{S}}$ are encoded by a T5 model, projected to length $F$, and added to their corresponding joint features across all frames. Finally, the block outputs enhanced data of shape $\mathbb{R}^{(N+1)\times J\times F}$.
@@
We extend transformers for graph-based learning by incorporating both graph topology and node interaction information through our Skeletal Attention mechanism. Inspired by discriminative works in the graphs domain (Ying et al., 2021b), AnyTop introduces a novel method for generative tasks, specifically tailored to the motion domain. We integrate graph properties directly into attention maps, enabling the structural characteristics of the graph to influence the learning process. Our work uses two types of node affinity, the topological distance, $\mathcal{D}_{\mathcal{S}}$, and relations, $\mathcal{R}_{\mathcal{S}}$, as detailed in section 3.1. We incorporate the graph information into the attention maps (Park et al., 2022), by learning distinct query and key embeddings for distances, denoted by $E^{\mathcal{D}}_{q}$, $E^{\mathcal{D}}_{k}\in\mathbb{R}^{d_{max}\times F}$, and embeddings for relation, denoted by $E^{\mathcal{R}}_{q}$, $E^{\mathcal{R}}_{k}\in\mathbb{R}^{6\times F}$, where $E^{(\cdot)}_{q}$ and $E^{(\cdot)}_{k}$ denote embeddings that relate to queries and keys, respectively, and $F$ is the latent feature size. These embeddings are used to form two new attention maps, $a^{\mathcal{D}}$ and $a^{\mathcal{R}}$ defined for a given pair of joints $i,j\in[J]$:
@@ 3.3. Training
We train AnyTop using minibatches sampled with a Balancing Sampler to address the imbalanced nature of the data (described in section 6.1) and mitigate the dominance of specific skeletons. To further enhance generalization, we apply skeletal augmentations to the data samples, including randomly removing 10% to 30% of the joints and adding new joints at the midpoint of existing edges. Further details on our data augmentation are provided in Appendix B .
@@
(4) $\mathcal{L}_{simple}=E_{t\sim[1,T]}\left\lVert\hat{X}_{0}-X_{0}\right\rVert^{2}_{2}.$
@@ 4.1. Latent Space Analysis
Our spatial and temporal correspondence results are illustrated in figs. 4 and 5 respectively, and in the supplementary video.
@@
We show that manifold features of semantically similar skeletal joints across different characters are close to each other. Our objective is to find the most similar joint in $\mathcal{S}^{ref}$ for each joint in $\mathcal{S}^{tgt}$. To achieve this, we extract DIFT features for both motions $X^{ref},X^{tgt}$ at diffusion step $t_{corr}=2$ and layer $l_{corr}=0$, average them along the temporal axis, and obtain a single feature vector per joint. Using cosine similarity, we detect the closest counterpart for each joint in $S_{tgt}$.
Temporal Correspondence
@@
We show that AnyTop can recognize pose-level similarities and identify analogous actions across different skeletons. This time, our objective is to find the most similar frame in $X^{ref}$ for each frame in $X^{tgt}$. To accomplish this goal, we extract DIFT features at diffusion step $t_{corr}=3$ and layer $l_{corr}=1$, and average them along the skeletal axis, resulting in a single feature vector per frame. We use cosine similarity to detect the closest counterpart for each frame in $X_{tgt}$.
@@ 4.2. Generalization Forms
extends to skeletons not encountered during training, and illustrated in fig. 8 and the video.
@@ 5. Applications
Temporal segmentation is the task of partitioning a temporal sequence into disjoint groups, where frames sharing similar characteristics are grouped. For a clean sample $X_{0}$, either generated or given, and skeleton $\mathcal{S}$, we extract DIFT features at diffusion step $t_{seg}\!=\!\!3$ and layer $l_{seg}\!=\!\!1$. The features are averaged along the joint dimension to produce a single feature vector per frame. We apply PCA for dimensionality reduction and then use K-means to cluster the frames into $k\!\!=\!\!3$ categories. Our results are visualized in fig. 9 and in the supp. video. This application reinforces section 4, showing that AnyTop’s latent features are effective frame descriptors. However, in section 4, frames are grouped by similarity to $X^{ref}$, while here they are grouped by similarity to each other.
@@
We demonstrate our method’s versatility through two motion editing applications: in-betweening for temporal manipulation and body-part editing for spatial modifications, both leveraging the same underlying approach. For in-betweening, the prefix and suffix of the motion are fixed, allowing the model to generate the middle. For body-part editing, we fix some of the joints and let the model generate the rest. Given a fixed subset (temporal or spatial) of the motion sequence tokens, we override the denoised $\hat{x}_{0}$ at each sampling iteration with the fixed motion part. This approach ensures fidelity to the fixed input while synthesizing the missing elements of the motion. Our results, in fig. 10 and the supp. video, show a smooth and natural transition between the given and the synthesized parts, and demonstrate that our model successfully generalizes techniques previously limited to human skeletons (Tevet et al., 2023) to accommodate diverse skeletal structures.
@@ 6. Experiments
Figure 8. Unseen-skeleton generalization Zero-shot inference of the cat (left) and komodo dragon (right) using AnyTop (top) and adapted MDM baseline (bottom). AnyTop’s generated motions maintain natural appearance while MDM’s generated motions are static and jittery.
[…]
@@ 6.1. Dataset and Preprocessing
The Truebones Zoo (Truebones Motions Animation Studios, 2022) dataset comprises motion captures featuring 70 diverse skeletons, including mammals, birds, insects, dinosaurs, fish, and snakes. The number of motions per skeleton ranges from 3 to 40, adding up to 1219 motions and 147,178 frames in total. The dataset includes variations in orientation, root definition, and scale. Additionally, the skeletons vary in joint order, naming conventions, and connectivity standards. To address these variations, we have performed comprehensive preprocessing of the data, including aligning all motions to the same orientation and average bone length, centering the first frame at the origin, and ensuring it is located on the ground. This process is described in details in Appendix B and the processed data will be made available.

### v2

@@ Abstract.
Generating motion for arbitrary skeletons is a longstanding challenge in computer graphics, remaining largely unexplored due to the scarcity of diverse datasets and the irregular nature of the data. In this work, we introduce AnyTop, a diffusion model that generates motions for diverse characters with distinct motion dynamics, using only their skeletal structure as input. Our work features a transformer-based denoising network, tailored for arbitrary skeleton learning, integrating topology information into the traditional attention mechanism. Additionally, by incorporating textual joint descriptions into the latent feature representation, AnyTop learns semantic correspondences between joints across diverse skeletons. Our evaluation demonstrates that AnyTop generalizes well, even with as few as three training examples per topology, and can produce motions for unseen skeletons as well. Furthermore, our model’s latent space is highly informative, enabling downstream tasks such as joint correspondence, temporal segmentation, and motion editing. Our webpage, https://anytop2025.github.io/Anytop-page, includes links to videos and code.
Animation, Motion synthesis, Deep Features, Computer Graphics.
copyright: noneccs: Computing methodologies Motion processingccs: Computing methodologies Computer graphicsccs: Computing methodologies Computer visionccs: Computing methodologies Machine learning approaches
Figure 1. AnyTop generates motions for diverse characters with distinct motion dynamics, using only their skeletal structure and joint names as input.
@@ 1. Introduction
AnyTop makes the following key contributions:
  • A new problem formulation for motion synthesis controlled by diverse skeletal structures.
  • A novel architecture generalizing to non-homeomorphic skeletons with significant diversity.
  • Comprehensive evaluation showing superior performance over baselines.
  • Careful analysis of learned features and generalization capabilities, establishing a foundation for interpretability and downstream tasks.
  • A unified version of the Truebones zoo dataset with a preprocessing framework.
2. Related Work
[…]
Skeletal variability in generative motion models
@@ 2. Related Work
We refer to four types of skeletal variability (Tab. 1). The naming draws from terminology in the graph domain, hence we interchangeably use the terms joint and vertex, as well as edge and bone. A single skeleton type refers to identical skeletons — that is, skeletons with the same vertices, connectivity, and edge lengths. Isomorphic skeletons correspond to isomorphic graphs, sharing vertices and edges but potentially differing in edge proportions. Homeomorphic skeletons may vary in structure, yet correspond to homeomorphic graphs, i.e., use topologies obtained from the same primal graph by subdivision of edges. Specifically, homeomorphic skeletons share the same number of kinematic chains and end-effectors. Finally, non-homeomorphic skeletons vary in their structure and have no common primal graph.
@@
Most motion generative methods focus on a single skeletal structure (Raab et al., 2023; Petrovich et al., 2021; Karunratanakul et al., 2023). Others train on isomorphic skeletons (Zhang et al., 2023b; Villegas et al., 2021), including works that use the SMPL (Loper et al., 2015) body model (Tripathi et al., 2025; Jang et al., 2024; Petrovich et al., 2022) and SMAL (Zuffi et al., 2017) body model or its derivatives (Yang et al., 2023; Rueegg et al., 2023). A smaller portion of generative works support homeomorphic skeletons (Lee et al., 2023; Zhang et al., 2024e; Cao and Yang, 2024; Studer et al., 2024; Ponton et al., 2024). Among these works, some (Aberman et al., 2020) require a designated encoder and decoder per skeleton, and some (Zhang et al., 2024b) offer a unified framework for all skeletons.
@@
Only a handful of works can handle non-homeomorphic skeletons. Martinelli et al. (2024) performs motion retargeting by learning a shared manifold for all skeletons, and decoding it to motions using learned skeleton-specific tokens. The learned tokens capture the skeletal information of characters in the dataset, limiting the model’s ability to generalize to skeletons unseen during training. Its results are shown exclusively on bipeds, leaving the applicability to other character families (e.g., quadrupeds, millipedes) unexplored. WalkTheDog (Li et al., 2024) uses a latent space that encodes motion phases and accommodates non-generative motion matching. A table clustering the methods mentioned above by supported skeletal variability can be found in Appendix A.
@@
A different class of generative models bypasses the handling of the skeletal structure by generating motion directly from point clouds (Mo et al., 2025), shape-handles (Zhang et al., 2023a) or meshes (Song et al., 2023; Ye et al., 2024; Muralikrishnan et al., 2024; Zhang et al., 2024d). These works demonstrate great flexibility in target character structure, but overlook the advantage of skeletons, which are more compact and semantically meaningful, easier to manipulate via rig-based animation, and compatible with physics engines (Tevet et al., 2024) and inverse kinematics systems. Some works (Wang et al., 2024) perform automatic rigging after the generation, but automatic rigging often necessitates manual adjustments.
@@ 3.1. Preliminaries
We represent motion as a 3D tensor $X\in\mathbb{R}^{N\times J\times D}$, where $N$ and $J$ are the maximum number of frames and joints across all motions in the dataset, and D is the number of motion features per joint. As motions vary in duration and skeletal structure, we pad the original number of frames and joints of each motion to match the maximum values $N$ and $J$, respectively. We adopt a redundant representation, where each joint $j$ (except the root) consists of its root-relative position $p_{j}\in\mathbb{R}^{3}$, 6D joint rotation $r_{j}\in\mathbb{R}^{6}$ (Zhou et al., 2018), linear velocity $v_{j}\in\mathbb{R}^{3}$, and foot contact label $fc_{j}\in\{0,1\}$. Altogether a joint is represented by $\{p_{j},r_{j},v_{j},fc_{j}\}\in\mathbb{R}^{13}$, hence $D=13$. Our representation is inspired by Guo et al. (2022); however, our approach maintains features at the joint level by representing each joint as a separate tensor, yielding $J$ tokens per frame. Conversely, Guo et al. concatenate features from all joints into one tensor, producing one token per frame.
@@ 3.2. Architecture
This block incorporates semantic information from the rest-pose $\mathcal{P}_{\mathcal{S}}$ and the joint descriptions $\mathcal{N}_{\mathcal{S}}$, into the noised sample $X_{t}$. It projects $\mathcal{P}_{\mathcal{S}}$ to latent feature size $F$ and concatenates it with $X_{t}$ along the temporal axis, effectively making it frame 0. The joint descriptions $\mathcal{N}_{\mathcal{S}}$ are encoded by a T5 model (Raffel et al., 2020), projected to length $F$, and added to their corresponding joint features across all frames. Finally, the block outputs the enhanced data of shape $\mathbb{R}^{(N+1)\times J\times F}$.
@@
We extend transformers for graph-based learning by incorporating both graph topology and node interaction information through our Skeletal Attention mechanism. Inspired by discriminative works in the graphs domain (Ying et al., 2021b), AnyTop introduces a novel method for generative tasks, specifically tailored to the motion domain. We integrate graph properties directly into attention maps, enabling the structural characteristics of the graph to influence the learning process. Our work uses two types of node affinity, the topological distance, $\mathcal{D}_{\mathcal{S}}$, and relations, $\mathcal{R}_{\mathcal{S}}$, as detailed in Sec. 3.1. We incorporate the graph information into the attention maps (Park et al., 2022), by learning distinct query and key embeddings for distances, denoted by $E^{\mathcal{D}}_{q}$, $E^{\mathcal{D}}_{k}\in\mathbb{R}^{(d_{max}+1)\times F}$, and embeddings for relation, denoted by $E^{\mathcal{R}}_{q}$, $E^{\mathcal{R}}_{k}\in\mathbb{R}^{6\times F}$, where $E^{(\cdot)}_{q}$ and $E^{(\cdot)}_{k}$ denote embeddings that relate to queries and keys, respectively, and $F$ is the latent feature size. These embeddings are used to form two new attention maps, $a^{\mathcal{D}}$ and $a^{\mathcal{R}}$ defined for a given pair of joints $i,j\in[J]$:
@@ 3.3. Training
We train AnyTop using minibatches sampled with a Balancing Sampler to mitigate data imbalance and the dominance of specific skeletons. Given $k$ skeletal types, with $n_{i}$ samples corresponding to skeletal type $i$, our balancing sampler assigns a sampling probability of $1/(n_{i}\cdot k)$ to each instance of type $i$. To further enhance generalization, we apply skeletal augmentations, including randomly removing 10% to 30% of the joints and adding new joints at the midpoint of existing edges. More details on data augmentation are provided in Appendix C.
@@
(4) $\mathcal{L}_{simple}=E_{t\sim[1,T]}\left\lVert AnyTop(X_{t},t,\mathcal{S})-X_{0}\right\rVert^{2}_{2}.$
@@ 4.1. Latent Space Analysis
Our spatial and temporal correspondence results are illustrated in Figs. 4 and 5, respectively, and in the supplementary video.
@@
We show that manifold features of semantically similar skeletal joints across different characters are close to each other (Fig. 4). Our objective is to find the most similar joint in $\mathcal{S}^{ref}$ for each joint in $\mathcal{S}^{tgt}$. To achieve this, we extract DIFT features for both motions $X^{ref},X^{tgt}$ at diffusion step $t_{corr}=2$ and layer $l_{corr}=0$, average them along the temporal axis, and obtain a single feature vector per joint. Using cosine similarity, we detect the closest counterpart for each joint in $S_{tgt}$.
Fig. 4 shows a clear alignment between corresponding body parts (e.g., limbs, spine, tail) across different animals, with only minor confusion related to left-right distinction. Left-right issues are common in feature extraction (Zhang et al., 2024c), to be addressed directly.
Temporal Correspondence
We show that AnyTop can recognize pose-level similarities and identify analogous actions across different skeletons (Fig. 5). This time, our objective is to find the most similar frame in $X^{ref}$ for each frame in $X^{tgt}$. To accomplish this goal, we extract DIFT features at diffusion step $t_{corr}=3$ and layer $l_{corr}=1$, and average them along the skeletal axis, resulting in a single feature vector per frame. We use cosine similarity to detect the closest counterpart for each frame in $X_{tgt}$.
@@ 4.2. Generalization Forms
extends to skeletons not encountered during training and illustrated in Fig. 8 and the video. We assess generalization to varying levels of skeletal out-of-distribution (OOD) using three insect characters: Crab (in-distribution), Centipede (moderately OOD), and Cobra (highly OOD). We quantify the OOD degree using Wasserstein distance on graph features. Each character is evaluated in both seen and unseen settings. Results, in Tab. 2, demonstrate robust performance on unseen skeletons that are in-distribution, with a gradual decline as the skeletons deviate from the training distribution. The decline is driven by divergence from the data distribution, featured by a drop in fidelity (coverage, intra-diversity difference) and a rise in local/inter-diversity.
Figure 8. Unseen-skeleton generalization Zero-shot inference of the cat (left) and komodo dragon (right) using adapted MDM baseline (top) and AnyTop (bottom). AnyTop’s generated motions maintain natural appearance while MDM’s generated motions are static and jittery.
[…]
Table 2. Unseen-skeleton generalization vs. OOD degree. Performance on unseen skeletons gradually drops as they deviate from data distribution.
Character OOD Score Coverage $\uparrow$ Seen / Unseen Local Div. $\uparrow$ Seen / Unseen Inter Div. $\uparrow$ Seen / Unseen Intra Div. Diff.$\downarrow$ Seen / Unseen
Crab 1.70 99.45 / 88.69 0.22 / 0.29 0.34 / 0.36 0.13 / 0.14
Centipede 2.26 83.55 / 43.90 0.28 / 0.29 0.41 / 0.16 0.12 / 0.15
Cobra 4.04 79.34 / 16.47 0.17 / 0.56 0.39 / 1.03 0.20 / 0.46
@@ 5. Applications
Temporal segmentation is the task of partitioning a temporal sequence into disjoint groups, where frames sharing similar characteristics are grouped. For a clean sample $X_{0}$, either generated or given, and skeleton $\mathcal{S}$, we extract DIFT features at diffusion step $t_{seg}\!=\!\!3$ and layer $l_{seg}\!=\!\!1$. The features are averaged along the joint dimension to produce a single feature vector per frame. We apply PCA for dimensionality reduction and then use K-means to cluster the frames into $k\!\!=\!\!3$ categories. Our results are visualized in Fig. 9 and in the supp. video. This application reinforces Sec. 4, showing that AnyTop’s latent features are effective frame descriptors. However, in Sec. 4, frames are grouped by similarity to $X^{ref}$, while here they are grouped by similarity to each other.
@@
We demonstrate our method’s versatility through two motion editing applications: in-betweening for temporal manipulation and body-part editing for spatial modifications, both leveraging the same underlying approach. For in-betweening, the prefix and suffix of the motion are fixed, allowing the model to generate the middle. For body-part editing, we fix some of the joints and let the model generate the rest. Given a fixed subset (temporal or spatial) of the motion sequence tokens, we override the denoised $\hat{x}_{0}$ at each sampling iteration with the fixed motion part. This approach ensures fidelity to the fixed For a clean sample $X_{0}$, either generated or given, and skeleton $\mathcal{S}$, we extract DIFT features at diffusion step $t_{seg}\!=\!\!3$ and layer $l_{seg}\!=\!\!1$. The features are averaged along the joint dimension to produce a single feature vector per frame. We apply PCA for dimensionality reduction and then use K-means to cluster the frames into $k\!\!=\!\!3$ categories. Our results are visualized in Fig. 9 and in the supp. video. while synthesizing the missing elements of the motion. Our results, in Fig. 10 and the supp. video, show a smooth and natural transition between the given and the synthesized parts, and demonstrate that our model successfully generalizes techniques previously limited to human skeletons (Tevet et al., 2023) to accommodate diverse skeletal structures.
@@ 6.1. Dataset and Preprocessing
The Truebones Zoo (Truebones Motions Animation Studios, 2022) dataset comprises motion captures featuring 70 diverse skeletons, including mammals, birds, insects, dinosaurs, fish, and snakes. The number of motions per skeleton ranges from 3 to 40, adding up to 1219 motions and 147,178 frames in total. The dataset includes variations in orientation, root definition, and scale. Additionally, the skeletons vary in joint order, naming conventions, and connectivity standards. To address these variations, we have performed comprehensive preprocessing of the data, including aligning all motions to the same orientation and average bone length, centering the first frame at the origin, and ensuring it is located on the ground. This process is described in detail in Appendix C, and the processed data will be made available.